# IllustrisTNG in the HSC-SSP: No Shortage of Thin Disk Galaxies in TNG50

### v1

@@ 5.1 The Significance of Equitable Comparisons
H22 first extracted the 3D axis lengths of the simulated galaxies defined as $\sqrt{\sum_{j}m_{j}r_{j,i}^{2}}/\sqrt{\sum_{j}m_{j}}$ from supplementary catalogs of the corresponding simulations (e.g., Genel et al., 2015, for TNG), where $m_{j}$ is the stellar mass of the $j$-th particle and $r_{j,i}$ is its $i$-th axis coordinate. They then projected an ellipsoid extended by such axis lengths onto a sky plane and built a distribution of projected axis ratios under the assumption of random lines of sight. This approach may not be appropriate for deriving comparable axis ratios to those obtained from the images. Firstly, the assumption of a single ellipsoid for their shape is oversimplified for disk galaxies. It blends all structural components together, therefore the spheroidal components, such as stellar bulges and halos, inevitably contribute to the 3D axis ratio calculations and then to the projected axis ratios at all inclinations. The fact that the 3D axis ratios are defined as the second moment of stellar mass distribution may exaggerate the contamination from the spheroidal components when compared with galaxy shapes measured from optical images, because (1) stellar bulges and halos are often fainter than disks at fixed stellar mass density due to their higher optical mass-to-light ratios and (2) they contribute substantially to the intrinsic flatness $q_{0}$—especially the stellar halo particles far away from the disk plane. In the EAGLE simulation, de Graaff et al. (2022) found that galaxy axis ratio distributions measured from projected stellar mass maps are consistent with those from their $r$-band images, which suggests that the former may not be the primary cause and the latter plays a more dominant role.
Our method of fitting an isophote in the outskirt of the projected galaxy images strives to avoid the center. Also, it does not reach the surface brightness level of stellar halos (Merritt et al., 2016). This approach provides reliable measurements of the projected shape of disks, except for edge-on galaxies where bulges inevitably come into play. It makes sense that the axis ratios obtained via parametric fitting (i.e., fitting 2D parametric surface brightness models to images) are broadly consistent with ours in terms of $f_{q\leq 0.4}$ (Figure 3), because both our method and parametric fitting is essentially an isophote-based method. Still, differences in the details remain. Parametric fitting assumes model isophotes to be concentric and to have the same shape, so the best-fit axis ratio is the average axis ratio of the overall flux distribution and would be biased by central components that have a different shape such as bulges. Our method, in contrast, is more robust against such bias.
@@ 5.2 The Low-mass Thicker Galaxies in TNG50
In addition, as highlighted by Xu et al. (2024, in preparation), axis ratio measurements for edge-on galaxies are affected by kinematically hot structures such as stellar bulges and other spheroidal components. Consequently, variations in structural compositions of galaxies may play an important role in shaping the dependence of axis ratios on stellar masses. The U-shape dependence of $q_{0}$ on stellar mass is reminiscent of that of $S/T$ on stellar mass in TNG, first reported by Tacchella et al. (2019) in TNG100 and reproduced in Figure 5 for TNG50. A similar trend was also found in the EAGLE simulation (Clauwens et al., 2018). The median $S/T$ reaches its minimum at $10.510<\log(M_{\star}/M_{\sun})<$ 10.5 in both TNG and EAGLE. The coincidence of the vertex stellar masses of the $q_{0}$$\log M_{\star}$ and $S/T$$\log M_{\star}$ relations and the stellar mass where early-type galaxies start to dominate in TNG (Huertas-Company et al., 2019) leads us to investigate whether the stellar mass dependence of galaxy structural compositions also contributes to the curvature in the $q_{0}$$\log M_{\star}$ relation of TNG50.
@@
In Figure 6, we compare the $C_{82}$$\log M_{\star}$ relation of the observed galaxies and the TNG50 galaxies. The $C_{82}$ is commonly recognized as an economical substitute for the Sersic (1968) index and bulge-to-total ratio (e.g., Gadotti, 2009). We find broad agreement between the two as well as some interesting differences. The median relations of the TNG50 galaxies exhibit a minimum at a similar stellar mass as the aforementioned U-shape relations, while the median relations of the HSC-SSP galaxies are monotonic. Because it is unclear how the spheroidal components quantitatively contribute to the $q_{0}$ measurements, the absolute difference of $C_{82}$ at fixed stellar masses between the observations and TNG50 is irrelevant for interpreting the shape of the $q_{0}$$\log M_{\star}$ relations. What matters is the relative difference of the trends with stellar mass between them.
@@
Towards lower stellar masses, the TNG50 relations show a less significant decrease in $C_{82}$ compared to the observational trends in all three bands. Towards higher stellar masses, the TNG50 galaxies exhibit a more prominent increase in $C_{82}$ than the observed ones in all three bands. These differences persist even after smoothing the trends by randomly perturbing the stellar masses of the simulated galaxies. This is also true for the comparison of concentrations and bulge statistics between TNG100 galaxies and real galaxies observed in the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; Chambers et al., 2016) $i$-band images at $z\approx 0.05$ (Rodriguez-Gomez et al., 2019). Assuming that more prominent spheroidal components at a given stellar mass lead to larger $q_{0}$ estimates of galaxies in the same sample, the difference in the $q_{0}$$\log M_{\star}$ relations is compatible with the more curved TNG50 $q_{0}$$\log M_{\star}$ relations than the HSC-SSP ones. It also explains why the U-shape is most noticeable in the reddest $y$ band, where the spheroidal components should be most prominent in flux (as seen in the elevation of $C_{82}$ from the $g$ to $y$ band; Figure 6), because they are mostly quenched structures in massive galaxies. However, we argue that this cannot be the dominant factor for the low-mass thicker galaxies in TNG50, otherwise we should observe a similar degree of mismatch between the observations and TNG50 at the high-mass end and should expect to observe thicker low-mass TNG50 galaxies in the $g$ and $i$ bands as well.
@@ 5.3 Difference in Galaxy Concentration between Observations and TNG50
The difference between observations and simulations at the low-mass end can be attributed to poorer numerical resolution. In addition, Rodriguez-Gomez et al. (2019) showed that different dust models yielded $\sim 0.1$ dex variations in the mock observed $i$-band concentration index of simulated galaxies at $\log(M_{\star}/M_{\sun})\approx 9.5$ (their Figure 6). Understanding the difference at the high-mass end requires more dedicated effort. Dust modeling in generating the synthetic images has a negligible effect since massive galaxies are mostly gas-poor (Rodriguez-Gomez et al., 2019). Guzmán-Ortega et al. (2023) found slightly higher merger rates in TNG50 than in KiDS galaxies using both Random Forrest and asymmetry-based classifications, which seems a plausible explanation for the higher concentrations in TNG50 galaxies. However, they also showed that merger and non-merger TNG50 galaxies have similar $r$-band concentrations. Moreover, classification of mergers suffers from incompleteness and impurity issues (e.g., Bickley et al., 2021; Bottrell et al., 2022). The uncertainties in stellar mass estimates of the observed galaxies could also contribute to the difference in the $C_{82}$$\log M_{\star}$ relations, as demonstrated in Rodriguez-Gomez et al. (2019) by randomly perturbing the stellar masses of the simulated galaxies and finding smoother trends. We strive to emulate such uncertainties following the approach of Rodriguez-Gomez et al. (2019) and find that the difference between TNG50 and observations remains significant (Figure 6). However, we caution that this approach does not account for the inclination dependent uncertainties in stellar mass estimates.
@@ 6 Summary
  • We find that low-mass galaxies ($\log(M_{\star}/M_{\sun})<9.5$) in TNG50 are thicker than their observed counterparts. We suggest that spurious dynamical heating of simulated stellar components by dark matter particles is the main cause, instead of unrealistic physics.
@@
Our study highlights the importance of conducting mock observations of simulations and employing consistent methods to extract measurements for comparison with observational results. The minor ($\lesssim 0.1\sigma$) difference in the overall axis ratio distributions between observations and the TNG50 simulation adds to the success of TNG in terms of reproducing realistic galaxy morphologies (e.g., Rodriguez-Gomez et al., 2019; Guzmán-Ortega et al., 2023; Eisert et al., 2024). Nevertheless, we also reveal differences in both the $q_{0}$$\log M_{\star}$ and the $C_{82}$$\log M_{\star}$ relations between the TNG50 simulation and the observation. This emphasizes the necessity to perform detailed comparisons of galaxy morphologies as a function of galaxy properties in addition to the overall distributions (Huertas-Company et al., 2019; Eisert et al., 2024). We discuss the possible origins of the difference in galaxy concentrations between the simulation and the observations. It will require further investigation to ascertain the primary cause.
@@
HMY was partially supported by JSPS KAKENHI Grant Number JP22K14072 and the Research Fund for International Young Scientists of NSFC (11950410492). We are deeply grateful for the opportunity to observe the Universe from Mauna Kea, which has always held a very significant cultural role within the indigenous Hawaiian community. The TNG50 simulation was run with compute time granted by the Gauss Centre for Supercomputing (GCS) under Large-Scale Projects GCS-DWAR on the GCS share of the supercomputer Hazel Hen at the High Performance Computing Center Stuttgart (HLRS). The Hyper Suprime-Cam (HSC) collaboration includes the astronomical communities of Japan and Taiwan, and Princeton University. The HSC instrumentation and software were developed by the National Astronomical Observatory of Japan (NAOJ), the Kavli Institute for the Physics and Mathematics of the Universe (Kavli IPMU), the University of Tokyo, the High Energy Accelerator Research Organization (KEK), the Academia Sinica Institute for Astronomy and Astrophysics in Taiwan (ASIAA), and Princeton University. Funding was contributed by the FIRST program from the Japanese Cabinet Office, the Ministry of Education, Culture, Sports, Science and Technology (MEXT), the Japan Society for the Promotion of Science (JSPS), Japan Science and Technology Agency (JST), the Toray Science Foundation, NAOJ, Kavli IPMU, KEK, ASIAA, and Princeton University. Funding for SDSS-III has been provided by the Alfred P. Sloan Foundation, the Participating Institutions, the National Science Foundation, and the U.S. Department of Energy Office of Science. The SDSS-III website is http://www.sdss3.org/. SDSS-III is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS-III Collaboration including the University of Arizona, the Brazilian Participation Group, Brookhaven National Laboratory, University of Cambridge, Carnegie Mellon University, University of Florida, the French Participation Group, the German Participation Group, Harvard University, the Instituto de Astrofisica de Canarias, the Michigan State/Notre Dame/JINA Participation Group, Johns Hopkins University, Lawrence Berkeley National Laboratory, Max Planck Institute for Astrophysics, Max Planck Institute for Extraterrestrial Physics, New Mexico State University, New York University, Ohio State University, Pennsylvania State University, University of Portsmouth, Princeton University, the Spanish Participation Group, University of Tokyo, University of Utah, Vanderbilt University, University of Virginia, University of Washington, and Yale University. GAMA is a joint European-Australasian project based around a spectroscopic campaign using the Anglo-Australian Telescope. The GAMA input catalog is based on data taken from the Sloan Digital Sky Survey and the UKIRT Infrared Deep Sky Survey. Complementary imaging of the GAMA regions is being obtained by a number of independent survey programs, including GALEX MIS, VST KiDS, VISTA VIKING, WISE, Herschel-ATLAS, GMRT, and ASKAP, providing UV to radio coverage. GAMA is funded by the STFC (UK), the ARC (Australia), the AAO, and the participating institutions. The GAMA website is http://www.gama-survey.org/.
@@ Appendix A The Effect of the Resolution Cut on Completeness of the Axis Ratio Distribution
Figure 7: The distributions of $y$-band axis ratios of the HSC-SSP sample (left) and the TNG50 sample (right) with resolution cut (open solid) and without resolution cut (filled dashed) in different stellar mass bins. Stellar mass increases from left to right and from top to bottom in the range $9\leq\log(M_{\star}/M_{\sun})\leq 11.5$. The parameter $f$ is the ratio of the number of galaxies post to pre-resolution cut in each stellar mass bin.

### v2

@@ 5.1 The Significance of Equitable Comparisons
H22 first extracted the 3D axis lengths of the simulated galaxies defined as $\sqrt{\sum_{j}m_{j}r_{j,i}^{2}}/\sqrt{\sum_{j}m_{j}}$ from supplementary catalogs of the corresponding simulations (e.g., Genel et al., 2015, for TNG), where $m_{j}$ is the stellar mass of the $j$th particle and $r_{j,i}$ is its $i$th axis coordinate. They then projected an ellipsoid extended by such axis lengths onto a sky plane and built a distribution of projected axis ratios under the assumption of random lines of sight. This approach may not be appropriate for deriving comparable axis ratios to those obtained from the images. Firstly, the assumption of a single ellipsoid for their shape is oversimplified for disk galaxies. It blends all structural components together, therefore the spheroidal components, such as stellar bulges and halos, inevitably contribute to the 3D axis ratio calculations and then to the projected axis ratios at all inclinations. The fact that the 3D axis ratios are defined as the second moment of stellar mass distribution may exaggerate the contamination from the spheroidal components when compared with galaxy shapes measured from optical images, because (1) stellar bulges and halos are often fainter than disks at fixed stellar mass density due to their higher optical mass-to-light ratios and (2) they contribute substantially to the intrinsic flatness $q_{0}$—especially the stellar halo particles far away from the disk plane. In the EAGLE simulation, de Graaff et al. (2022) found that galaxy axis ratio distributions measured from projected stellar mass maps are consistent with those from their $r$-band images, which suggests that the former may not be the primary cause and the latter plays a more dominant role.
Our method of fitting an isophote in the outskirt of the projected galaxy images strives to avoid the center. Also, it does not reach the surface brightness level of stellar halos (Merritt et al., 2016). This approach provides reliable measurements of the projected shape of disks, except for edge-on galaxies where bulges inevitably come into play. It makes sense that the axis ratios obtained via parametric fitting (i.e., fitting 2D parametric surface brightness models to images) are broadly consistent with ours in terms of $f_{q\leq 0.4}$ (Figure 3), because both our method and parametric fitting are essentially isophote-based methods. Still, differences in the details remain. Parametric fitting assumes model isophotes to be concentric and to have the same shape, so the best-fit axis ratio is the average axis ratio of the overall flux distribution and would be biased by central components that have a different shape such as bulges. Our method, in contrast, is more robust against such bias.
@@ 5.2 The Low-mass Thicker Galaxies in TNG50
In addition, as highlighted by D. Xu et al. (2024, in preparation), axis ratio measurements for edge-on galaxies are affected by kinematically hot structures such as stellar bulges and other spheroidal components. Consequently, variations in structural compositions of galaxies may play an important role in shaping the dependence of axis ratios on stellar masses. The U-shape dependence of $q_{0}$ on stellar mass is reminiscent of that of S/T on stellar mass in TNG, first reported by Tacchella et al. (2019) in TNG100 and reproduced in Figure 5 for TNG50. A similar trend was also found in the EAGLE simulation (Clauwens et al., 2018). The median S/T reaches its minimum at $10.510<\log(M_{\star}/M_{\sun})<$ 10.5 in both TNG and EAGLE. The coincidence of the vertex stellar masses of the $q_{0}$$\log M_{\star}$ and S/T–$\log M_{\star}$ relations and the stellar mass where early-type galaxies start to dominate in TNG (Huertas-Company et al., 2019) leads us to investigate whether the stellar mass dependence of galaxy structural compositions also contributes to the curvature in the $q_{0}$$\log M_{\star}$ relation of TNG50.
@@
In Figure 6, we compare the $C_{82}$$\log M_{\star}$ relation of the observed galaxies and the TNG50 galaxies. The $C_{82}$ is commonly recognized as an economical substitute for the Sérsic (1968) index and bulge-to-total ratio (e.g., Gadotti, 2009). We find broad agreement between the two as well as some interesting differences. The median relations of the TNG50 galaxies exhibit a minimum at a similar stellar mass as the aforementioned U-shape relations, while the median relations of the HSC-SSP galaxies are monotonic. Because it is unclear how the spheroidal components quantitatively contribute to the $q_{0}$ measurements, the absolute difference of $C_{82}$ at fixed stellar masses between the observations and TNG50 is irrelevant for interpreting the shape of the $q_{0}$$\log M_{\star}$ relations. What matters is the relative difference of the trends with stellar mass between them.
@@
Toward lower stellar masses, the TNG50 relations show a less significant decrease in $C_{82}$ compared to the observational trends in all three bands. Toward higher stellar masses, the TNG50 galaxies exhibit a more prominent increase in $C_{82}$ than the observed ones in all three bands. These differences persist even after smoothing the trends by randomly perturbing the stellar masses of the simulated galaxies. This is also true for the comparison of concentrations and bulge statistics between TNG100 galaxies and real galaxies observed in the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; (Chambers et al., 2016) $i$-band images at $z\approx 0.05$ (Rodriguez-Gomez et al., 2019). Assuming that more prominent spheroidal components at a given stellar mass lead to larger $q_{0}$ estimates of galaxies in the same sample, the difference in the $q_{0}$$\log M_{\star}$ relations is compatible with the more curved TNG50 $q_{0}$$\log M_{\star}$ relations than the HSC-SSP ones. It also explains why the U-shape is most noticeable in the reddest $y$ band, where the spheroidal components should be most prominent in flux (as seen in the elevation of $C_{82}$ from the $g$ to $y$ band; Figure 6), because they are mostly quenched structures in massive galaxies. However, we argue that this cannot be the dominant factor for the low-mass thicker galaxies in TNG50, otherwise we should observe a similar degree of mismatch between the observations and TNG50 at the high-mass end and should expect to observe thicker low-mass TNG50 galaxies in the $g$ and $i$ bands as well.
@@ 5.3 Difference in Galaxy Concentration between Observations and TNG50
The difference between observations and simulations at the low-mass end can be attributed to poorer numerical resolution. In addition, Rodriguez-Gomez et al. (2019) showed that different dust models yielded $\sim 0.1$ dex variations in the mock observed $i$-band concentration index of simulated galaxies at $\log(M_{\star}/M_{\sun})\approx 9.5$ (their Figure 6). Understanding the difference at the high-mass end requires more dedicated effort. Dust modeling in generating the synthetic images has a negligible effect since massive galaxies are mostly gas-poor (Rodriguez-Gomez et al., 2019). Guzmán-Ortega et al. (2023) found slightly higher merger rates in TNG50 than in KiDS galaxies using both random forest and asymmetry-based classifications, which seems a plausible explanation for the higher concentrations in TNG50 galaxies. However, they also showed that merger and non-merger TNG50 galaxies have similar $r$-band concentrations. Moreover, classification of mergers suffers from incompleteness and impurity issues (e.g., Bickley et al., 2021; Bottrell et al., 2022). The uncertainties in stellar mass estimates of the observed galaxies could also contribute to the difference in the $C_{82}$$\log M_{\star}$ relations (Rodriguez-Gomez et al., 2019). We strive to emulate such uncertainties by randomly perturbing the stellar masses of the simulated galaxies and find that the difference between TNG50 and the observations remains significant despite the smoother trends in TNG50 (Figure 6). However, we caution that this approach does not account for the inclination dependent uncertainties in stellar mass estimates.
@@ 6 Summary
  • We find that low-mass galaxies ($\log(M_{\star}/M_{\sun})\lesssim 9.5$) in TNG50 are thicker than their observed counterparts. We suggest that spurious dynamical heating of simulated stellar components by dark matter particles is the main cause, instead of unrealistic physics.
@@
Our study highlights the importance of conducting mock observations of simulations and employing consistent methods to extract measurements for comparison with observational results. The minor ($\lesssim 0.1\sigma$) difference in the overall axis ratio distributions between observations and the TNG50 simulation adds to the success of TNG in terms of reproducing realistic galaxy morphologies (e.g., Rodriguez-Gomez et al., 2019; Guzmán-Ortega et al., 2023; Eisert et al., 2024). Our results call into question the previous claims that $\Lambda$CDM produced too few thin galaxies compared to observations (e.g., Kormendy et al., 2010; van de Sande et al., 2019; H22). Nevertheless, we also reveal differences in both the $q_{0}$$\log M_{\star}$ and the $C_{82}$$\log M_{\star}$ relations between the TNG50 simulation and the observation. This emphasizes the necessity to perform detailed comparisons of galaxy morphologies as a function of galaxy properties in addition to the overall distributions (Huertas-Company et al., 2019; Eisert et al., 2024). We discuss the possible origins of the difference in galaxy concentrations between the simulation and the observations. It will require further investigation to ascertain the primary cause.
@@
H.G. thanks the support from the Pan-STARRS project, which is supported by NASA 80NSSC21K1572 and by the State of Hawaii. C.B. gratefully acknowledges the support of the Forrest Research Foundation. H.M.Y. was partially supported by JSPS KAKENHI Grant Number JP22K14072 and the Research Fund for International Young Scientists of NSFC (11950410492). We are deeply grateful for the opportunity to observe the Universe from Mauna Kea, which has always held a very significant cultural role within the indigenous Hawaiian community. The TNG50 simulation was run with compute time granted by the Gauss Centre for Supercomputing (GCS) under Large-Scale Projects GCS-DWAR on the GCS share of the supercomputer Hazel Hen at the High Performance Computing Center Stuttgart (HLRS). The Hyper Suprime-Cam (HSC) collaboration includes the astronomical communities of Japan and Taiwan, and Princeton University. The HSC instrumentation and software were developed by the National Astronomical Observatory of Japan (NAOJ), the Kavli Institute for the Physics and Mathematics of the Universe (Kavli IPMU), the University of Tokyo, the High Energy Accelerator Research Organization (KEK), the Academia Sinica Institute for Astronomy and Astrophysics in Taiwan (ASIAA), and Princeton University. Funding was contributed by the FIRST program from the Japanese Cabinet Office, the Ministry of Education, Culture, Sports, Science and Technology (MEXT), the Japan Society for the Promotion of Science (JSPS), Japan Science and Technology Agency (JST), the Toray Science Foundation, NAOJ, Kavli IPMU, KEK, ASIAA, and Princeton University. Funding for SDSS-III has been provided by the Alfred P. Sloan Foundation, the Participating Institutions, the National Science Foundation, and the U.S. Department of Energy Office of Science. The SDSS-III website is http://www.sdss3.org/. SDSS-III is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS-III Collaboration including the University of Arizona, the Brazilian Participation Group, Brookhaven National Laboratory, University of Cambridge, Carnegie Mellon University, University of Florida, the French Participation Group, the German Participation Group, Harvard University, the Instituto de Astrofisica de Canarias, the Michigan State/Notre Dame/JINA Participation Group, Johns Hopkins University, Lawrence Berkeley National Laboratory, Max Planck Institute for Astrophysics, Max Planck Institute for Extraterrestrial Physics, New Mexico State University, New York University, Ohio State University, Pennsylvania State University, University of Portsmouth, Princeton University, the Spanish Participation Group, University of Tokyo, University of Utah, Vanderbilt University, University of Virginia, University of Washington, and Yale University. GAMA is a joint European-Australasian project based around a spectroscopic campaign using the Anglo-Australian Telescope. The GAMA input catalog is based on data taken from the Sloan Digital Sky Survey and the UKIRT Infrared Deep Sky Survey. Complementary imaging of the GAMA regions is being obtained by a number of independent survey programs, including GALEX MIS, VST KiDS, VISTA VIKING, WISE, Herschel-ATLAS, GMRT, and ASKAP, providing UV to radio coverage. GAMA is funded by the STFC (UK), the ARC (Australia), the AAO, and the participating institutions. The GAMA website is http://www.gama-survey.org/.
@@ Appendix A The Effect of the Resolution Cut on Completeness of the Axis Ratio Distribution
Figure 7: The distributions of $y$-band axis ratios of the HSC-SSP sample (left) and the TNG50 sample (right) with resolution cut (open solid) and without resolution cut (filled dashed) in different stellar mass bins. Stellar mass increases from left to right and from top to bottom in the range $9\leq\log(M_{\star}/M_{\sun})\leq 11.5$. The parameter $f$ is the ratio of the number of galaxies after the resolution cut to before the resolution cut in each stellar mass bin.